Larry's Auto Body Repair Shop had revenues that averaged $70,000 per week in April and $60,000 per week in May. During both months, the shop employed six full-time (40 hours/week) workers. In April the firm also had four part-time workers working 25 hours per week but in May there were only 2 part-time workers and they only worked 10 hours per week. What is the percentage change in labor productivity from April to May for Larry's Auto Body Repair?

a.

+8.97

b.

+19.23

c.

-1.76

d.

+12.09

e.

-15.84
Answer #1

ANSWER d: +12.09%

EXPLANATION:

For April per week data,

Number of labor hours = (6 x 40) + (4 x 25) = 340

labour productivity = $70000/340 = $205.88 per hour

for May per week data,

Number of labor hours = (6 x 40) + (2 x 10) = 260

labour productivity = $60000/260 = $230.77 per hour

percentage change in labor productivity = ($230.77 - $205.88)/$205.88

= +12.09%